# Thinking in Cocktail Party: Chain-of-Thought and Reinforcement Learning for Target Speaker Automatic Speech Recognition

### v1

@@ Abstract
Target Speaker Automatic Speech Recognition (TS-ASR) aims to transcribe the speech of a specified target speaker from multi-speaker mixtures in cocktail party scenarios. Recent advancement of Large Audio-Language Models (LALMs) has already brought some new insights to TS-ASR. However, significant room for optimization remains for the TS-ASR task within the LALMs architecture. While Chain of Thoughts (CoT) and Reinforcement Learning (RL) have proven effective in certain speech tasks, TS-ASR, which requires the model to deeply comprehend speech signals, differentiate various speakers, and handle overlapping utterances is particularly well-suited to a reasoning-guided approach. Therefore, we propose a novel framework that incorporates CoT and RL training into TS-ASR for performance improvement. A novel CoT dataset of TS-ASR is constructed, and the TS-ASR model is first trained on regular data and then fine-tuned on CoT data. Finally, the model is further trained with RL using selected data to enhance generalized reasoning capabilities. Experiment results demonstrate a significant improvement of TS-ASR performance with CoT and RL training, establishing a state-of-the-art performance compared with previous works of TS-ASR on comparable datasets.
Index Terms—  Target Speaker Automatic Speech Recognition, Chain-of-Thought, Reinforcement Learning
@@ 1 Introduction
Large Language Models (LLMs), which originally developed for text understanding and generation, have achieved significant advances by scaling the parameters and the data exponentially [1, 2]. In the field of speech and audio processing, Large Audio-Language Models (LALMs) have developed rapidly in recent years, expanding the capabilities of LLMs to speech-related tasks through pre-trained speech encoders [3, 4]. When dealing with Automatic Speech Recognition (ASR) tasks, for example, prevailing architectures of LLM-based ASR typically employ a pre-trained speech encoder to transform acoustic signals into feature representations, which are then injected into LLMs as prompts for end-to-end ASR training [5, 6]. Furthermore, LALMs architecture holds the potential to extend capabilities to more complex acoustic environments, such as TS-ASR in cocktail party scenarios, which remains a challenging research frontier.
@@
Given a reference speech from the target speaker, TS-ASR is capable of selectively transcribing the speech of the target speaker while suppressing interference from other speakers [7]. Some previous works designed a neural network architecture to jointly train Target Speaker Extraction (TSE) and ASR in order to transcribe the speech of the target speaker [8, 9]. There were also some works integrating speaker embeddings into the training of conventional end-to-end ASR models to directly train the TS-ASR model [10, 11, 12]. With the recent advancement of LALMs, MT-LLM [13], which utilized both an ASR encoder and a speaker encoder to extract features and then fed them into the LLM backbone to get TS-ASR output, has also demonstrated a good performance on the TS-ASR task. Although the development of LALMs has already brought new insights into TS-ASR, the optimization potential remains far from exhausted, as TS-ASR is a complex task that requires deep audio comprehension to achieve higher accuracy.
@@
The breakthrough of DeepSeek-R1 [14] shows that combining CoT with RL significantly improves reasoning in complex tasks. This approach has also proven to be effective on many speech tasks, such as Audio Question and Answering (AQA) [15, 16] and ASR [17]. TS-ASR is a complex task that requires models to deeply understand speech and distinguish between different speakers, which involves logical reasoning to determine the number of speakers, identify the target speaker, and select the relevant speech segments for recognition. Therefore, we believe that CoT and RL are able to improve the performance of TS-ASR task by guiding the model to produce explicit intermediate reasoning steps before arriving at the final answer.
@@
In this work, we introduce CoT and RL into the TS-ASR task to enhance reasoning capabilities in cocktail party scenarios. First, a TS-ASR BASE model is trained based on the LALMs architecture. Then, a novel method for constructing CoT training data is proposed, in which key information such as speaker count, overlap duration, speaker gender, and similarity to the reference is extracted from mixed speech. These attributes are subsequently structured through logical organization to form reasoning data for the CoT training, which is fine-tuned on the TS-ASR BASE model. Finally, based on the model trained on CoT data, part of the incorrectly predicted training data are randomly selected to do the further RL training. Experiment results show that our proposed method achieves an averaged 8.33% Word Error Rate (WER) of LibriSpeech, Libri2Mix and Libri3Mix. As far as we know, our method achieves a State-Of-The-Art (SOTA) performance compared with previous works of TS-ASR [9, 12, 13] on comparable datasets.
Fig. 1: Overview of the proposed method.
@@
  • We pioneer the integration of CoT and RL training into the TS-ASR task, establishing a novel framework that enhances reasoning capabilities in cocktail party scenarios.
  • We construct a novel CoT training dataset for TS-ASR by extracting task-relevant information in logic from the mixed speech. This CoT dataset will be open-sourced to support further research at https://github.com/Ease-Z/TS-ASR-CoT
  • Our approach demonstrates a significant improvement of TS-ASR performance with CoT and RL training, establishing a SOTA performance compared with previous works of TS-ASR on comparable datasets.
@@ 2.1.1 Model structure
The proposed framework of the TS-ASR model is shown in Figure 1 (a). It consists of speech encoder, linear adapter, text tokenizer, and an LLM backbone network. We select a 3s segment from the target speaker utterance as reference, concatenate it with the mixed input (separated by a 3s silence), and feed the concatenated audio into the speech encoder. The pre-trained speech encoder Data2Vec2 [18] converts the input speech into frame-level embedding vectors that contain semantic and speaker information. To achieve modality alignment, a linear adapter maps the output from the speech encoder to the feature space of LLM. Meanwhile, the text tokenizer transforms the textual prompt into text embeddings. The speech and text embeddings are then combined and fed into the LLM. After fine-tuning, the model can focus on the target speaker within the mixed speech and generate accurate transcriptions.
In our framework, reference speech is concatenated with mixed speech as an audio prompt [19]. This approach effectively leverages the inherent capabilities of LLMs in generating targeted responses from prompts for TS-ASR. Moreover, the architecture employs only the Data2Vec2 speech encoder without introducing a separate speaker encoder. As speaker and semantic information are both inherently captured by the Self-Supervised Learning (SSL) model [20], their fusion is naturally achieved within the encoder.
Fig. 2: Sample of CoT data for single-speaker and 2-speaker mixed speech. Text in different colors represents different types of information.
@@ 2.1.2 Training pipeline
Our TS-ASR training follows a three-stage paradigm as shown in Figure 1 (b):
TS-ASR BASE A base TS-ASR model is trained via full parameter Supervised Fine-Tuning (SFT) on the initialized LALM, where all components are jointly updated, including the pre-trained LLM, the pre-trained speech encoder, and the randomly initialized adapter. This step aims to align multi-modal information and enable the model to perform TS-ASR. It effectively addresses the cold start problem and provides a foundation for subsequent CoT training.
CoT Training The TS-ASR BASE model is fine-tuned with constructed CoT data to guide step-by-step reasoning to improve TS-ASR performance. To focus model capacity on challenging cases, a random reasoning strategy is applied, providing full CoT procedure for hard samples while keeping empty <think>for easy samples.
RL Training Furthermore, based on the CoT trained model, RL refines the reasoning steps and improves the final performance using Generalized Reward Policy Optimization (GRPO) [21] with WER and CoT format rewards. In order to further focus learning on challenging cases, the RL training set is constructed by selecting a subset of constructed CoT data with wrong format or ASR transcription.
@@ 2.2.1 CoT data construction
Our CoT dataset is constructed based on LibriSpeech [22] dataset, which contains gender information, speaker information, and ASR transcription for each single utterance. As shown in Figure 2, since the reasoning structure required for TS-ASR is relatively fixed, a standardized data template is designed that includes five components. First, audio information describes the composition of the input speech, the number of speakers, and the total duration. Second, reference speech primarily records the gender of the target speaker. Third, the corresponding speech of each speaker is described by timestamps, gender, and a similarity level to the reference. Since the mixed speech is generated from single-speaker utterances, these information can be obtained from the source single-speaker speech. Fourth, the target speaker is then inferred based on the above information. In single-speaker scenarios, identification is straightforward, while in multi-speaker scenarios, gender and similarity level are jointly considered to reach a decision. Finally, the answer is the transcription corresponding to the target speaker.
To compute similarity levels, we use CAM++ [23] to extract speaker embeddings and calculate the cosine similarity scores between embeddings of each source speech and the reference speech. Then, these continuous scores in $[0,1]$ are mapped to five discrete levels $[1,5]$ via uniform quantization. This discretization prevents the model from focusing on insignificant numerical differences, thereby improving training stability and convergence efficiency.
@@ 2.2.2 Training with CoT data
The constructed CoT data are employed to further fine-tune the TS-ASR BASE model via SFT. Since LLMs often struggle with complex tasks, CoT guides them with explicit reasoning steps like humans. In this stage, the model is trained to analyze the speech content, estimate the number of speakers, describe the attributes of each speaker, and reason step by step to identify the target speaker. Finally, the model generates transcription from the identified target speaker.
Moreover, we introduce a random reasoning strategy for CoT training data. CoT data with the complete thinking is provided for hard samples, which is defined as those mispredicted by the BASE model. While the <think>content is left empty with 50% probability for easy samples that are correctly predicted. This strategy encourages the model to focus more on reasoning for challenging samples while allowing it to output answers directly for easier cases.
@@ 2.3.1 Data selection
Since RL fine-tuning can only use a subset of the training data, a data selection strategy is designed for TS-ASR, focusing on samples most beneficial for performance improvement. Specifically, the full training set is decoded using the model trained by CoT data, selecting all samples with incorrectly predicted CoT format and a random subset of samples with correct format but wrong ASR prediction. This approach is motivated by the fact that format errors are more fundamental and impactful than content errors, and including them in training helps the model further refine its reasoning process.
2.3.2 Generalized reward policy optimization
@@
The reward $r$ for GRPO training is the combination of WER reward $r_{\text{WER}}$ and CoT format reward $r_{\text{format}}$:
@@
where $F$ is the format: <think></think><answer></answer><think>...</think><answer>...</answer> and $y$ is the output string. The format reward encourages the model to perform the full reasoning process and ensures correct reasoning structure.
@@ 3.1 Dataset
Based on the approach in LibriMix [24], this study generates mixed speech datasets of two speakers (Libri2Mix) and three speakers (Libri3Mix) using the 960-hour LibriSpeech [22] single-speaker corpus [^1] https://github.com/JorisCos/LibriMix. Following the data scale in  [13], we create mixed samples for each source utterance, alternately selecting each speaker in the mixture as the target. Reference speech segments are randomly selected from the single utterances of the same speaker. The duration of the Libri2Mix and Libri3Mix datasets is 1920 and 2880 hours, respectively. When combined with the original 960 hours of single-speaker data, the total training set reaches about 5760 hours. The test set is generated from LibriSpeech test-clean, with Libri2Mix and Libri3Mix constructed using the mixture and reference speech provided by the open-source SpeakerBeam toolkit [^2] https://github.com/BUTSpeechFIT/speakerbeam/tree/main/egs.

### v2

@@ Abstract
Target Speaker Automatic Speech Recognition (TS-ASR) aims to transcribe the speech of a specified target speaker from multi-speaker mixtures in cocktail party scenarios. Recent advancement of Large Audio-Language Models (LALMs) has already brought some new insights to TS-ASR. However, significant room for optimization remains for the TS-ASR task within the LALMs architecture. While Chain of Thoughts (CoT) and Reinforcement Learning (RL) have proven effective in certain speech tasks, TS-ASR, which requires the model to deeply comprehend speech signals, differentiate various speakers, and handle overlapping utterances is particularly well-suited to a reasoning-guided approach. Therefore, we propose a novel framework that incorporates CoT and RL training into TS-ASR for performance improvement. A novel CoT dataset of TS-ASR is constructed, and the TS-ASR model is first trained on regular data and then fine-tuned on CoT data. Finally, the model is further trained with RL using selected data to enhance generalized reasoning capabilities. Experiment results show a significant improvement of TS-ASR performance with CoT and RL training, which demonstrates the effectiveness of the proposed CoT and RL training methods adapted for the TS-ASR task.
Machine Learning, ICML
@@ 1 Introduction
Large Language Models (LLMs), which originally developed for text understanding and generation, have achieved significant advances by scaling the parameters and the data exponentially (Achiam et al., 2023; Touvron et al., 2023). In the field of speech and audio processing, Large Audio-Language Models (LALMs) have developed rapidly in recent years, expanding the capabilities of LLMs to speech-related tasks through pre-trained speech encoders (Huang et al., 2024; Zhang et al., 2023a). When dealing with Automatic Speech Recognition (ASR) tasks, for example, prevailing architectures of LLM-based ASR typically employ a pre-trained speech encoder to transform acoustic signals into feature representations, which are then injected into LLMs as prompts for end-to-end ASR training (Yu et al., 2024; Bai et al., 2024). Furthermore, LALMs architecture holds the potential to extend capabilities to more complex acoustic environments, such as Target-Speaker ASR (TS-ASR) in cocktail party scenarios, which remains a challenging research frontier.
Speech recognition in multi-talker scenario aims to transcribe speech signals containing overlapping utterances from multiple speakers into text. Current tasks in this scenario mainly fall into two categories: Multi-Talker ASR (MT-ASR) and Target-Speaker ASR (TS-ASR). MT-ASR (Moriya et al., 2025; Shi et al., 2024) seeks to transcribe all speakers present in a mixed audio signal. The dominant approach is Serialized Output Training (SOT)(Kanda et al., 2020) which outputs transcriptions sequentially, typically separated by a special token, following a “first-in, first-out” strategy. While this simplifies the recognition process, it does not associate each transcription with a specific speaker, making it more suitable for applications like meeting transcription.
Figure 1: The overview of TS-ASR framework based on LALM architecture.
@@
In contrast, many applications such as smart home assistants, and conversational agents require accurately recognizing speech of a specific user in noisy, multi-speaker environments. To address these requirements, TS-ASR has been widely studied. Given a reference speech of the target speaker, TS-ASR selectively transcribes the speech of the target speaker while suppressing interference from others (Kanda et al., 2019). The TS-ASR task relies on additional reference information and is sensitive to speaker verification errors, making it a challenge task in multi-talker scenario. Some previous works designed a neural network architecture to jointly train Target Speaker Extraction (TSE) and ASR in order to transcribe the speech of the target speaker (Zhang et al., 2023b; Huang et al., 2023). There were also some works integrating speaker embeddings into the training of conventional end-to-end ASR models to directly train the TS-ASR model (Guo et al., 2025b; Ma et al., 2024; Maeda et al., 2025). With the recent advancement of LALMs, MT-LLM (Meng et al., 2025), which utilized both an ASR encoder and a speaker encoder to extract features and then fed them into the LLM backbone to get TS-ASR output, has also demonstrated a good performance on the TS-ASR task. Although the development of LALMs has already brought new insights into TS-ASR, the optimization potential remains far from exhausted, as TS-ASR is a complex task that requires deep audio comprehension to achieve higher accuracy.
@@
The breakthrough of DeepSeek-R1 (Guo et al., 2025a) shows that combining CoT with RL significantly improves reasoning in complex tasks. This approach has also proven to be effective on many speech tasks, such as Audio Question and Answering (AQA) (Rouditchenko et al., 2025; Ma et al., 2025b) and ASR (Shivakumar et al., 2025). TS-ASR is a complex task that requires models to deeply understand speech and distinguish between different speakers, which involves logical reasoning to determine the number of speakers, identify the target speaker, and select the relevant speech segments for recognition. Therefore, we believe that CoT and RL are able to improve the performance of TS-ASR task by guiding the model to produce explicit intermediate reasoning steps before arriving at the final answer.
@@
In this work, we design a set of CoT and RL training method adapted for the TS-ASR task to enhance reasoning capabilities in cocktail party scenarios. First, a TS-ASR BASE model is trained based on the LALMs architecture. Then, a novel method for constructing CoT training data is proposed, in which key information such as speaker count, overlap duration, speaker gender, and similarity to the reference is extracted from mixed speech. These attributes are subsequently structured through logical organization to form reasoning data for the CoT training, which is fine-tuned on the TS-ASR BASE model. Finally, based on the model trained on CoT data, RL training is further conducted with selected data to enhance the generalization capabilities of the model. Experiment results show a significant improvement of TS-ASR performance with proposed CoT training method, and a further improvement is achieved after conducting the proposed RL training method. This demonstrates the effectiveness of the proposed CoT and RL training method designed for the TS-ASR task.
Figure 2: Training pipeline of the proposed method.
@@
  • This is the first work to propose that logical reasoning benefits the TS-ASR task, and a set of CoT and RL training methods adapted for the TS-ASR task are thereby proposed.
  • We construct a novel CoT training dataset for TS-ASR by extracting task-relevant information in logic from the mixed speech. This CoT dataset will be open-sourced to support further research at https://anonymous.4open.science/r/TS-ASR-CoT-RL-anonymous-1DFF
  • Our approach shows a significant improvement of TS-ASR performance with CoT and RL training, which demonstrates the effectiveness of the proposed CoT and RL training methods adapted for the TS-ASR task, establishing a novel framework that enhances reasoning capabilities in cocktail party scenarios.
@@ 2.1.1 Model structure
The proposed framework of the TS-ASR model is shown in Figure 1. It consists of speech encoder, linear adapter, text tokenizer, and an LLM backbone network. We select a 3s segment from the target speaker utterance as reference, concatenate it with the mixed input (separated by a 3s silence), and feed the concatenated audio into the speech encoder. The pre-trained speech encoder Data2Vec2 (Baevski et al., 2023) converts the input speech into frame-level embedding vectors that contain semantic and speaker information. To achieve modality alignment, a linear adapter maps the output from the speech encoder to the feature space of LLM. Meanwhile, the text tokenizer transforms the textual prompt into text embeddings. The speech and text embeddings are then combined and fed into the LLM. After fine-tuning, the model can focus on the target speaker within the mixed speech and generate accurate transcriptions.
In our framework, reference speech is concatenated with mixed speech as an audio prompt (Cheng and Nguyen, 2025). This approach effectively leverages the inherent capabilities of LLMs in generating targeted responses from prompts for TS-ASR. Moreover, the architecture employs only the Data2Vec2 speech encoder without introducing a separate speaker encoder. As speaker and semantic information are both inherently captured by the Self-Supervised Learning (SSL) model (Chen et al., 2022), their fusion is naturally achieved within the encoder.
@@ 2.1.2 Training pipeline
Our TS-ASR training follows a three-stage paradigm as shown in Figure 2:
TS-ASR BASE A base TS-ASR model is trained via full parameter Supervised Fine-Tuning (SFT) on the initialized LALM, where all components are jointly updated, including the pre-trained LLM, the pre-trained speech encoder, and the randomly initialized adapter. This step aims to align multi-modal information and enable the model to perform TS-ASR. This step leverages TS-ASR data that pairs multi-talker speech with target speaker references, effectively addresses the cold start problem and provides a foundation for subsequent CoT training.
Figure 3: Sample of CoT data for single-speaker, 2-speaker and 3-speaker mixed speech. Text in different colors represents different types of information.
CoT Training The TS-ASR BASE model is fine-tuned with constructed CoT data to guide step-by-step reasoning. This teaches the model to first analyze the input speech by describing the reference speech and the attributes of each speaker, then reason step by step to identify the target speaker, and finally generate the transcription. The core idea is that, for complex tasks, LLMs often struggle to produce correct answers directly. The CoT method addresses this limitation by explicitly providing reasoning steps, guiding the model to think step by step like humans and thereby improving its reasoning capability.
RL Training Furthermore, based on the CoT trained model, RL refines the reasoning steps and improves the final performance using Group Relative Policy Optimization (GRPO) (Shao et al., 2024) with Word Error Rate (WER) and CoT format rewards. In order to further focus learning on challenging cases, the RL training set is constructed by selecting a ”Hard Data” subset of the original CoT data, which consists of examples where the CoT-trained model predicted either incorrectly CoT format or erroneous ASR transcriptions.
@@ 2.2.1 CoT data construction
Our CoT dataset is constructed based on LibriSpeech (Panayotov et al., 2015) dataset, which contains gender information, speaker information, and ASR transcription for each single utterance. As shown in Figure 3, since the reasoning structure required for TS-ASR is relatively fixed, a standardized data template is designed that includes five components. First, audio information describes the composition of the input speech, the number of speakers, and the total duration. Second, reference speech primarily records the gender of the target speaker. Third, the corresponding speech of each speaker is described by timestamps, gender, and a similarity level to the reference. Since the mixed speech is generated from single-speaker utterances, these information can be obtained from the source single-speaker speech. Fourth, the target speaker is then inferred based on the above information. In single-speaker scenarios, identification is straightforward, while in multi-speaker scenarios, gender and similarity level are jointly considered to reach a decision. Finally, the answer is the transcription corresponding to the target speaker.
To compute similarity levels, we use CAM++ (Wang et al., 2023) to extract speaker embeddings and calculate the cosine similarity scores between embeddings of each source speech and the reference speech. Then, these continuous scores in [1,1][-1,1] are mapped to five discrete levels $[1,5]$, defined as:
$l=\begin{cases}1,&\text{if }s<0,\\ 1+\lfloor 5\times s\rfloor,&\text{if }0\leq s<1,\\ 5,&\text{if }s=1.\end{cases}$ (1)
where $s$ denotes the similarity score and $l$ denotes the similarity level. Specifically, the continuous score $s$ is multiplied by 5, rounded down to the nearest integer, and then shifted by an offset of 1 to map the result into the range from 1 to 5. For the few cases where $s<0$, the score is clamped to the minimum level of 1. This discretization prevents the model from focusing on insignificant numerical differences, thereby improving training stability and convergence efficiency.
@@ 2.2.2 Training with CoT data
Since LLMs often struggle with complex tasks, CoT guides them with explicit reasoning steps like humans (Wei et al., 2022). In this stage, the model is trained to analyze the speech content, estimate the number of speakers and describe the attributes of each speaker. Based on the structured analysis, the model reasons step by step to determine which speaker is the target. Finally, the model generates transcription from the identified target speaker. This staged reasoning not only improves model performance but also enhances the interpretability and reliability of its predictions.
@@ 2.3.1 Data selection
High quality data benefits the RL training(Latif et al., 2023), therefore a data selection strategy is designed for the RL training of TS-ASR, focusing on samples most beneficial for performance improvement. As shown in Figure 2, inference is performed on the full training set using the model trained by CoT data, selecting all samples with incorrectly predicted CoT format and a random subset of samples with correct format but wrong ASR results. ASR errors serve as a direct indicator of task difficulty, while format errors are fundamental but important for format maintaining. Including both types of error data in RL training helps the model further refine its reasoning capabilities within the specified format. With this data selection strategy, RL training can focus on challenging samples, improving both discrimination and generalization capabilities on challenging cases.
2.3.2 Group Relative Policy Optimization
@@
In GRPO, for each input $x$, a group of $G$ outputs is generated and compared within the group based on task-specific rewards $r_{k}$. The normalized advantage $\hat{A}_{k}$ of the $k$-th response $y_{k}$ is:
$\hat{A}_{k}=\frac{r_{k}-\mathrm{mean}({\{r_{1},...,r_{G}\}})}{\mathrm{std}(\{r_{1},...,r_{G}\})}$ (2)
where $r_{k}$ is the reward of the response $k$.
The final GRPO objective is defined as:
$\begin{split}\mathcal{J}_{\mathrm{GRPO}}(\theta)=\frac{1}{G}\sum_{k=1}^{G}\frac{1}{|\boldsymbol{y}_{k}|}\sum_{t=1}^{|\boldsymbol{y}_{k}|}&\min\big\{\rho_{k}(\theta)\cdot\hat{A}_{k},\\ &\mathrm{clip}\big(\rho_{k}(\theta),1-\epsilon,1+\epsilon\big)\cdot\hat{A}_{k}\big\}\end{split}$ (3)
where the clip function limits the relative probability ratio $\rho_{k}(\theta)$ to a range defined by the hyperparameter $\epsilon$.
$\rho_{k}(\theta)=\frac{\pi_{\theta}(y_{k}\mid x)}{\pi_{\theta_{\mathrm{old}}}(y_{k}\mid x)}$ (4)
Here, $\pi_{\theta}$ and $\pi_{\theta_{\mathrm{old}}}$ are the current and old policy models, respectively.
2.3.3 GRPO for TS-ASR
In TS-ASR, the reward $r$ for GRPO training is the combination of WER reward $r_{\text{WER}}$ and CoT format reward $r_{\text{format}}$:
@@ 2.3.3 GRPO for TS-ASR
where $F$ is the format: <think></think><answer></answer><think>...</think><answer>...</answer> and $y$ is the output string. The format reward is designed to address cases where the model occasionally incompletely generates the required <think>and <answer>tags. It checks the completeness of these structural elements and encourages the model to perform the full reasoning process and ensures correct reasoning structure.
@@ 3.1 Dataset
Based on the approach in LibriMix (Cosentino et al., 2020), this study generates mixed speech datasets of two speakers (Libri2Mix) and three speakers (Libri3Mix) using the 960-hour LibriSpeech (Panayotov et al., 2015) single-speaker corpus [^1] https://github.com/JorisCos/LibriMix. Following the data scale in  (Meng et al., 2025), we create mixed samples for each source utterance, alternately selecting each speaker in the mixture as the target. Reference speech segments are randomly selected from the single utterances of the same speaker. The duration of the Libri2Mix and Libri3Mix datasets is 1920 and 2880 hours, respectively. When combined with the original 960 hours of single-speaker data, the total training set reaches about 5760 hours. The test set is generated from LibriSpeech test-clean, with Libri2Mix and Libri3Mix constructed using the mixture and reference speech provided by the open-source SpeakerBeam toolkit [^2] https://github.com/BUTSpeechFIT/speakerbeam/tree/main/egs.